OpenStudy (anonymous):

Matrices question: X-2y+z=-4 3x-4y+2z=3 4x-6y+3z=6 I did this question and got x, y, z = 11, 37/4,-7/2 But the answer given with the worksheet says no solution for this question so I was wondering if someone can tell me if I did the question incorrectly or the worksheet is wrong. Thanks! :)

OpenStudy (lyrae):

\[\left[\begin{matrix}1 & -2 & 1 & | & -4 \\ 3 & -4 & 2 & | & 3 \\ 4 & -6 & 3 & | & 6 \end{matrix}\right] \rightarrow \left[\begin{matrix}1 & -2 & 1 & | & -4 \\ 0 & 2 & -1 & | & 15 \\ 0 & 2 & -1 & | & 22 \end{matrix}\right] \rightarrow \left[\begin{matrix}1 & -2 & 1 & | & -4 \\ 0 & 2 & -1 & | & 15 \\ 0 & 0 & 0 & | & 7 \end{matrix}\right]\]This results in \[0 = 7\]which is impossible.

OpenStudy (lyrae):

So there's no answer.

OpenStudy (anonymous):

Thanks!

OpenStudy (lyrae):

Yw :)